### v1

# Assessing the Feasibility of Internet-Sourced Video for Automatic Cattle Lameness Detection

Abstract
Cattle lameness is often caused by hoof injuries or interdigital dermatitis, leads to pain and significantly impacts essential physiological activities such as walking, feeding, and drinking. This study presents a deep learning-based model for detecting cattle lameness, sickness, or gait abnormalities using publicly available video data. The dataset consists of 50 unique videos from 40 individual cattle, recorded from various angles in both indoor and outdoor environments. Half of the dataset represents naturally walking (normal/non-lame) cattle, while the other half consists of cattle exhibiting gait abnormalities (lame). To enhance model robustness and generalizability, data augmentation was applied to the training data. The pre-processed videos were then classified using two deep learning models: ConvLSTM2D and 3D CNN. A comparative analysis of the results demonstrates strong classification performance. Specifically, the 3D CNN model achieved a video-level classification accuracy of 90%, with precision, recall, and f1-score of 90.9%, 90.9%, and 90.91% respectively. The ConvLSTM2D model exhibited a slightly lower accuracy of 85%. This study highlights the effectiveness of directly applying classification models to learn spatiotemporal features from video data, offering an alternative to traditional multi-stage approaches that typically involve object detection, pose estimation, and feature extraction. Besides, the findings demonstrate that the proposed deep learning models, particularly the 3D CNN, effectively classify and detect lameness in cattle while simplifying the processing pipeline.
Index Terms:
Cattle, Lameness detection, Deep-learning, Computer vision techniques, Image processing
I Introduction
Figure 1: Overall workflow of the cattle lameness classification model
Cattle farms significantly contribute to global food security by providing meat and dairy products, with continuously increasing demand. While large-scale farming is now common, it presents operational challenges and necessitates continuous monitoring of animal health and welfare, as this directly impacts farm productivity [1, 2]. However, one of the leading health issues in cattle is lameness, a common and painful disorder frequently caused by gait-related injuries. Research indicates that approximately one in four cattle is affected by lameness at any given time, which reduces farm productivity and plays a substantial role in increasing meat and dairy prices [3, 5, 4]. A common practice about identifying lame cattle is herd locomotion scoring which requires visual observation of each animal individually. This process is cost and time consuming since they tend to increase human commissioning [6, 7]. Therefore, it is highly necessary to deploy advanced automatic technologies for identification of such health difficulties.
In recent years, AI through ML and data analysis have come up with smart solutions to serve large-scale cattle farming in different ways, e.g., by early detection of diseases and infections, by improving the production characteristics, etc [8]. AI uses data collected from livestock farms using various advanced tools including cameras, drones, robotics, and relevant IoT devices [9]. Although manual video cameras have been commonly used for a long time to surveillance the activities of animals from farmers’ desks, nowadays, computer vision and pattern recognition techniques have added new opportunities using the video data to automatically monitor animals’ health and welfare 24/7. Whether using a manual or computer vision (CV) approach, lameness is detected by observing behavioral changes in cattle—how they react and respond during regular activities in their living environment. Advanced deep learning-oriented AI models analyze image or video data from different perspectives, depending on the animal’s behavior (e.g., eating, walking, and general movement) [10]. Abnormal behavior is often readily apparent when cattle walk, as lameness is closely associated with painful gait and foot disease. The automation primarily focuses on detecting deviations between normal and abnormal behavior, classifying them to make decisions and further analysis.
However, the impact of such development in livestock farming is enormous, it reduces the labor costs, helps to identify health issues immediately, and increases production efficiency. Therefore, researchers are continuously working on the innovative smart farming solution with image processing and CV.
Fuentes et al. [9] used two Hikvision surveillance cameras to collect video data for their study. The cameras captured video from two different angles and extracted frames based on animal activity. They then used these frames to create a dataset, which was employed to train a model that effectively monitored the cows and performed statistical analysis to capture behavioral changes. McDonagh et al. [7] introduced a neural network model for behavior prediction by monitoring 46 individual pregnant cows. They specially considered the 10 hours prior video data to calving and focused on seven different behaviors. The model showed successful classification of the behavior over 80% instances and was able to predict calving before several hours with 83% accuracy. Fernández-Carrión et al. [8] proposed a deep learning model for the early detection of disease in Eurasian wild boar by monitoring their motion. They found a negative correlation between the animals’ movement and fever caused by the disease; the movement of infected animals was significantly reduced compared to that of healthy ones. In another study, Qiao et al. [10] proposed a combined model of Convolutional 3D and Convolutional Long Short-Term Memory (LSTM), which achieved higher accuracy compared to other state-of-the-art models in classifying five behaviors. Maria et al. [11]. focused on swine health, specifically identifying respiratory diseases. They considered eye temperature, heart rate, and respiration rate by extracting information from thermal infrared images. Their results demonstrated the successful implementation of CV in the early detection of respiratory diseases by considering physiological changes in pigs.
In another article, Barney et al. [13]. proposed a modified Mask-RCNN model for real-time automated lameness detection with high accuracy. In [14], the authors used Mask-RCNN and YOLO, both object detection algorithms. After detecting the cattle, they assigned a local ID to track each animal, then extracted features, and applied machine learning (ML) algorithms. Similarly, Russello et al. [5] presented a deep learning model to detect lameness by extracting nine annotations from video data. Wu et al. [12] introduced an LSTM model for large-scale farming to identify lame cows automatically. They used 50 videos, with a lame-to-non-lame sample ratio of 20:30. From these videos, they randomly selected 5,000 frames for their experiment, focusing solely on the position and movement of the cows’ legs. Their LSTM-based classification model achieved 98.57% accuracy, outperforming three traditional machine learning algorithms using a 50:50 train-test split.
However, the literature shows that cattle lameness was investigated in previous studies using surveillance video data collected from farms, with cameras installed at various angles. Additionally, most of those studies focused on small samples, and the investigations were conducted within limited and specific contexts. This study aimed to investigate lameness detection in large cattle farms, where hundreds of cattle need to be managed simultaneously. The research was conducted in two parts: video data collection and the development of a classification model. To achieve this, YouTube videos containing both lame and non-lame cattle were used. The labeling of normal and lame cattle was performed based on the title and description of each video. Since online videos were captured from diverse environments and locations, dataset diversity and robustness were ensured. To enhance the dataset size, vertical augmentation was applied to the training data. Moreover, to build the classification model, two deep learning classifiers, ConvLSTM2D and 3D CNN were utilized. The models were evaluated using multiple performance metrics. The classification results indicate successful implementation of online data with the proposed models, achieving comparable or superior accuracy compared to state-of-the-art models.
II Materials and Methods
TABLE I: Basic information of the dataset used in this study with the most relevant datasets in the literature
Study Dataset Size Test Environment Focused Activity
Train Frame Test Frame
1 1376 345 Within the barn Intensive farming
2 1000 500 Passing alley Intensive farming
3 490 210 Walked along a narrow path Intensive farming
4 690 341 Barn to milking parlor pathway Intensive farming
5 600 120 The corridor after the milking parlor Intensive farming
This study 3000 1000 Indoor and outdoor environment Semi-Intensive farming
Figure 1 illustrates the deep learning-based cattle lameness detection pipeline. The process begins with video data, from which frames are extracted during the data preprocessing stage. These extracted frames undergo data augmentation to enhance model generalization, producing a diverse training dataset. The model construction phase involves training two deep learning architectures: 3D CNN and ConvLSTM2D, denoted as $f_{\text{CNN3D}}(X)$ and $f_{\text{ConvLSTM2D}}(X)$, respectively. Finally, the models are evaluated using test data in the cattle classification stage to assess their performance.
II-A Data
The data used in previous relevant research papers were collected following a similar approach. First, a specific cattle farm, including cows, was selected. Then, different types of video cameras were installed in convenient locations on the grid. The cameras captured videos of the cattle from desired angles (e.g., walking, eating, and moving) for a specific amount of time. These videos were processed, and only the relevant parts were used for further analysis. Additionally, many studies performed locomotion scoring (a manual lameness detection approach) to classify cattle as lame or non-lame. Based on this classification, unique identification was assigned to the cattle, and tracking of both classes was maintained. Later, deep learning models were trained and assessed using the data.
II-A1 Data collection
The data collection process for this study was quite different from that of previous studies. We used an online platform to collect freely accessible cattle walking videos, specifically from YouTube. Publicly available data are commonly used in many research fields; for example, one of our previous articles [15] demonstrated that YouTube videos are frequently used in deepfake video detection research. In order to create an inclusive classification model, we created a diverse dataset. Therefore, the following features were incorporated with the video: there were different cattle in different environments, including intensive, semi-intensive, and extensive cattle farming; the cattle walked in different directions, including left to right and right to left; and the videos were captured from various views, including front, side, and back. For this study, a total of 50 videos featuring 40 cattle were considered, with an average duration of five seconds per video. We ensured an equal class distribution of the data, meaning that 25 lame and 25 non-lame cattle videos were included.
Table I provides an overview of the dataset used in this research alongside datasets from previous cattle lameness detection studies. The table compares these studies based on the number of training and testing frames, the test environment, and the focused activity. As shown in the last row, this study employs a comparatively larger dataset, consisting of 3,000 training and 1,000 testing frames. Previous studies focused on intensive farming environments, where cattle were primarily managed in indoor grazing conditions and data was collected from specific, limited locations such as within a barn, a passing alley, a narrow path, and corridors related to milking parlors. In contrast, this study focuses on a semi-intensive farming setting and incorporates both indoor and outdoor environments. This approach captures a wider range of cattle movement behaviors, lighting conditions, and surface variations, potentially leading to a more generalizable model for cattle lameness detection.
II-A2 Data pre-processing
Preprocessing in this investigation was conducted in several distinct steps, including frame extraction and data augmentation. Initially, frames were extracted at the standard rate of 30 frames per second. However, due to the high volume of frames, only 50 images were randomly selected from each training video, resulting in a total of 1500 images. These images were then cropped to a 608×608 pixel ratio. However, data augmentation is a widely recognized method for addressing limited data issues and helps to create robust and sustainable classification models. Numerous data augmentation techniques are available, as demonstrated by Jia et al. [3], who employed nine different methods, such as Mixup, Cutout, CutMix, etc. This study considered only image flipping as an augmentation technique, applied solely to the training data. All 30 training videos (1500 frames) were virtually flipped, generating a total of 3000 images.
II-B Model Construction
Algorithm 1 Pseudo-code of the proposed classification model
[…]
2:Trained 3DCNN OR ConvLSTM2D model; Performance metrics
3:1. Data Preparation:
4:for $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{test}}$ do
5:     $F_{i}\leftarrow\text{Pad/Truncate}(\text{ResizeFrames}(\text{ExtractFrames}(V_{i},\text{max\_frames}=50),S_{\text{target}}),N_{\text{frames}})$
6:     $(X,y)\leftarrow(X\cup\{F_{i}\},y\cup\{l_{i}\})$ if $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}$ else $(X_{\text{test}},y_{\text{test}})$
7:end for
8:Augment $X_{\text{train}},y_{\text{train}}$ with horizontal flipping (probability = $p_{\text{aug}}$).
9:Normalize and encode labels for $X_{\text{train}},y_{\text{train}},X_{\text{test}},y_{\text{test}}$.
10:2. Model Training:
11:$M\leftarrow\text{Create 3DCNN OR ConvLSTM2D Model}(N_{\text{frames}},S_{\text{target}})$
[…]
13:3. Model Evaluation:
14:$\hat{y}\leftarrow M(X_{\text{test}})$
15:metrics, confusion matrix $\leftarrow\text{Evaluate}(y_{\text{test}},\hat{y})$
16:return $M$, metrics, confusion matrix
The algorithm 1 outlines the pseudocode of the classfication model. It begins with data preparation by processing the training and testing datasets ($\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{test}}$), extracting frames from each video, resizing them to a target size ($S_{\text{target}}$), and ensuring each video has $N_{\text{frames}}$ frames through padding or truncation, creating input data ($X$) and labels ($y$). Training data is augmented with random horizontal flipping (probability $p_{\text{aug}}$). Pixel values are normalized, and behavior labels (“normal” and “lame”) are encoded numerically. Next, model training involves creating a 3D CNN or ConvLSTM2D model, configured using $N_{\text{frames}}$ and $S_{\text{target}}$, and training it on the prepared data ($X_{\text{train}}$, $y_{\text{train}}$) using a batch size ($B$) and epochs ($E$). The model learns patterns indicative of different behavior classes. Finally, model evaluation uses the trained 3D CNN model to predict behavior labels ($\hat{y}$) for the testing dataset ($X_{\text{test}}$), comparing them to true labels ($y_{\text{test}}$) to assess performance. Metrics like accuracy are calculated, and a confusion matrix is generated. The trained model, performance metrics, and confusion matrix are returned as output.
II-B1 3D CNN
CNNs are widely used deep learning models for image processing tasks. 3D CNNs are an extension of these and can handle 3D data containing spatiotemporal information, such as videos. They utilize 3D convolutional kernels, which enable them to focus on spatial features extracted from input data and distinguish themselves from traditional 2D CNNs. Because of their ability to learn both spatial and temporal features, this study employed a basic 3D CNN model to classify cattle videos, where movement patterns are assessed over time. The model extracts features by processing a sequence of video frames as input data through two convolutional layers (each followed by a max-pooling layer). The extracted features are then transformed into a one-dimensional vector by a flattening layer which is further used in two fully connected dense layers with ReLU activation. Two dropout layers are used to prevent overfitting, Finally, a softmax output layer provides the probability distribution over the defined classes.
II-B2 ConvLSTM2D
ConvLSTM2D is a deep learning architecture that integrates the strengths of spatial feature extraction and temporal dependency modeling through the combination of CNNs and LSTM networks, respectively. This makes it particularly effective for understanding patterns and movements in sequential data, such as videos. The ConvLSTM2D model employed in this study begins with a ConvLSTM2D layer. The model also incorporates two MaxPooling3D layers, one Conv3D layer, and dropout layers. Similar to the previous model, a dense layer with softmax activation is used to classify cattle into ‘normal’ and ‘lame’ categories.
II-C Model Evaluation
The four performance evaluation measures (accuracy, precision, recall, and F1-score) used in this study are associated with the confusion matrix; therefore, both are presented gradually. The values of all evaluation criteria range from 0% to 100%, where a higher value indicates stronger performance. In the mathematical framework for model evaluation, $n$ represents the total instances, $y_{i}$ denotes the true label for instance $i$, and $\hat{y}_{i}$ signifies the model’s predicted label. The indicator function $I(\text{condition})$ returns 1 if the condition is true, otherwise 0. Summation, denoted by $\sum_{i=1}^{n}$, calculates the cumulative value across all instances. These components enable quantitative assessment by comparing predicted ($\hat{y}_{i}$) and true ($y_{i}$) labels, facilitating calculation of the evaluation metrics, providing a comprehensive evaluation of model performance.
$Accuracy=\frac{1}{n}\sum_{i=1}^{n}I(y_{i}=\hat{y}_{i})$ (1)
$Precision=\frac{\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=1)}{\sum_{i=1}^{n}I(\hat{y}_{i}=1)}$ (2)
$Recall=\frac{\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=1)}{\sum_{i=1}^{n}I(y_{i}=1)}$ (3)
$F1-Score=\frac{2\times Precision\times Recall}{Precision+Recall}$ (4)
III Results
(a) 3DCNN
(b) ConvLSTM2D
Figure 2: Confusion Matrix of both models under the same test data
TABLE II: Lameness classification results of proposed 3D CNN and ConvLSTM2D
Model Accuracy Precision Recall F1-score
3DCNN 90 90.9 90.9 90.91
ConvLSTM2D 85 90 81.82 85.71
The dataset utilized in this study comprised 50 videos, with 50% depicting defective cattle and 50% depicting non-defective cattle. Of these 50 videos, 30 were used for model training, and 20 were used for validating the 3D CNN and ConvLSTM2D models. This project was developed using the Python programming language and implemented within the Google Colab Pro environment. Data preprocessing tasks were handled by OpenCV and MoviePy, supplemented by Keras’ ImageDataGenerator for data augmentation. Furthermore, TensorFlow, NumPy, Pandas, Scikit-learn, and Matplotlib were employed for model construction, training, performance evaluation, and visualization. The classification results achieved by the models were calculated and presented below.
The models generate predictions for each individual frame within each video in the test data. It is assumed that if the model correctly predicts the majority of frames in a video, then the entire video is considered correctly classified, despite the presence of a few incorrect frame-level predictions. The confusion matrices for both models are presented in Figure 2. As shown in Figures 2a and 2b, a small number of misclassifications were observed, including 1 false positives and 1 false negative for 3D CNN, and 1 false positives and 2 false negatives for ConvLSTM2D, out of the total 20 samples (10 normal and 10 lame).
The 3D CNN model demonstrated notable performance in classifying the cattle data. Table II indicates an accuracy of 90%, along with strong average precision, recall, and F1-scores. On the other hand, ConvLSTM2D achieved comparatively lower performance, with accuracy, precision, recall, and F1-score values of 85%, 90%, 81.82%, and 85.71%, respectively. Overall, the confusion matrices and classification reports highlight the models’ effectiveness in classifying healthy and defective cattle.
IV Discussion
TABLE III: Performance comparison of this study with existing cattle lameness classification models
Study Classifier Accuracy Precision Recall F1-score
[4] Logistic Regression 81.18 75 47.06 57.83
[10] C3D-ConvLSTM 90.32 90.62 88.76 89.68
[5] SVM 80.07 - - 78.70
[14] AdaBoost 77.9 - - -
[16] DT 90.18 - - -
[17] CNN 90.21 - - -
[18] SVM 83.58 86.04 88.43 87.21
[19] KNN 73.33 89.08 75.28 81.6
This study 3DCNN 90 90.9 90.9 90.91
  • - No specific data was found
Based on available information, all previous research on cattle lameness, sickness, or gait abnormality detection using computer vision technology has been conducted with very specific and limited data types. Typically, these studies focused on a single cattle barn, where video cameras were set up to monitor the cattle. The collected video data were then used for further analysis, including object detection, pose estimation, feature extraction, and classification. In contrast, this study utilized cattle videos from online platforms, which encompass multi-dimensional features. This departure from previous studies allowed for the development of a more universal and comprehensive lameness detection model that is suitable for large-scale cattle farms. The results of this study also validate the applicability of the classification model.
Table III presents a performance comparison between eight prior studies and the proposed model. The table evaluates models based on accuracy, precision, recall, and f1-score. The final row of the table displays the results of this study, which employed a 3D CNN model, while the preceding rows summarize the findings from other relevant studies that utilized various machine learning and deep learning classifiers. The model used in this study achieved a notable accuracy of 90%. More significantly, it demonstrated the highest precision among the studies where precision values were available, reaching 90.9%. This indicates that the model exhibits a strong ability to minimize false positives, correctly identifying positive cases with high confidence. Additionally, the 3D CNN model also yielded a robust F1-score of 90.91%. When compared to other approaches, the 3D CNN model in this study demonstrated competitive performance, whereas study [10] employed C3D-ConvLSTM achieved a slightly higher accuracy of 90.32% but a lower precision of 90.62%.
The results indicate that deep learning models generally outperform traditional machine learning approaches, such as SVM, Logistic Regression, and KNN. Furthermore, the findings of this study demonstrate the effectiveness of 3D CNNs for the video classification, as they efficiently capture both spatial and temporal information and offer well-balanced performance across multiple evaluation metrics.
IV-A Limitation and future plan
While this study demonstrates the potential of deep learning for cattle lameness detection, certain limitations may affect its broader applicability and real-world deployment, potentially impacting the generalizability of the findings to broader contexts. The following points outline these limitations:
  • All of the videos were sourced from YouTube. The cattle abnormality was determined based on the video titles and descriptions; however, it was not possible to individually verify the class and level of abnormality.
  • Although the simplicity of the proposed methodology (which directly employs classification algorithms) enables successful implementation with the video data, it is also necessary to experiment with intermediate processing steps such as object detection and pose estimation.
  • While this study utilized a greater number of training and testing frames compared to many similar studies, the total number of video samples remains relatively limited, which may affect the model’s generalizability and robustness.
To address the limitations outlined in this study and further enhance the applicability of our approach, future work should focus on expanding the dataset with a greater variety of video sources, including direct recordings from cattle farms, which would ensure a more robust representation of real-world scenarios, encompassing diverse cattle breeds, environmental conditions, and lameness severities, as well as addressing potential biases in online video content. In addition, while this study initially utilizes a direct classifier, future research will explore the impact of feature engineering through object detection, pose estimation, and feature extraction to assess the dataset and model performance. Overall, this study is an initial initiative of a benchmark cattle lameness detection dataset, future work will be introducing a complete benchmark of the approach.
V Conclusions
This paper presents a cattle lameness video dataset and deep learning-based methods for effectively distinguishing between lame and non-lame cattle from video data. Compared to existing datasets used for cattle lameness detection, this study collected data entirely from online sources, providing more comprehensive information across diverse environments, scenes, and activities. Additionally, two deep learning models, 3D CNN and ConvLSTM2D, were used to classify normal and lame cattle. From the dataset, 3,000 images were extracted for training the models, while 1,000 images were used for validation. The experimental results demonstrated improved or at least comparable performance to existing cattle lameness classification models. In conclusion, this study serves as a reference for detecting cattle lameness or behavior without relying on traditional multi-step approaches.