### v2

# Assessing the Feasibility of Internet-Sourced Video for Automatic Cattle Lameness Detection

Abstract
Cattle lameness is often caused by hoof injuries or interdigital dermatitis, leads to pain and significantly impacts essential physiological activities such as walking, feeding, and drinking. This study presents a deep learning-based model for detecting cattle lameness, sickness, or gait abnormalities using publicly available video data. The dataset consists of 50 unique videos from 42 individual cattle, recorded from various angles in both indoor and outdoor environments. Half of the dataset represents naturally walking (normal/non-lame) cattle, while the other half consists of cattle exhibiting gait abnormalities (lame). To enhance model robustness and generalizability, data augmentation was applied to the training data. The pre-processed videos were then classified using two modified deep learning models: ConvLSTM2D and 3D CNN. A comparative analysis of the results demonstrates strong classification performance. Specifically, the 3D CNN model achieved a video-level classification accuracy of 90%, with precision, recall, and f1-score of 90.9%, 90.9%, and 90.91% respectively. The ConvLSTM2D model exhibited a slightly lower accuracy of 85%. This study highlights the effectiveness of directly applying classification models to learn spatiotemporal features from video data, offering an alternative to traditional multi-stage approaches that typically involve object detection, pose estimation, and feature extraction. Besides, the findings demonstrate that the proposed deep learning models, particularly the 3D CNN, effectively classify and detect lameness in cattle while simplifying the processing pipeline.
Index Terms:
Cattle, Lameness detection, Deep-learning, Computer vision techniques, Image processing.
I Introduction
Figure 1: Overall workflow of the cattle lameness classification model
Cattle farms significantly contribute to global food security by providing meat and dairy products, with continuously increasing demand. While large-scale farming is now common, it presents operational challenges and necessitates continuous monitoring of animal health and welfare, as this directly impacts farm productivity [1, 2]. However, one of the leading health issues in cattle is lameness, a common and painful disorder frequently caused by gait-related injuries. Research indicates that approximately one in four cattle is affected by lameness at any given time, which reduces farm productivity and plays a substantial role in increasing meat and dairy prices [3, 5, 4]. A common practice about identifying lame cattle is herd locomotion scoring which requires visual observation of each animal individually [6, 7, 8, 9]. This process is cost and time consuming since they tend to increase human commissioning. Therefore, it is highly necessary to deploy advanced automatic technologies for identification of such health difficulties.
In recent years, AI through ML and data analysis have come up with smart solutions to serve large-scale cattle farming in different ways, e.g., by early detection of diseases and infections, by improving the production characteristics, etc [10]. AI uses data collected from livestock farms using various advanced tools including cameras, drones, robotics, and relevant IoT devices [14]. Although manual video cameras have been commonly used for a long time to surveillance the activities of animals from farmers’ desks [11, 12], nowadays, computer vision and pattern recognition techniques have added new opportunities using the video data to automatically monitor animals’ health and welfare 24/7 [13]. Whether using a manual or computer vision (CV) approach, lameness is detected by observing behavioral changes in cattle—how they react and respond during regular activities in their living environment. Advanced deep learning-oriented AI models analyze image or video data from different perspectives, depending on the animal’s behavior (e.g., eating, walking, and general movement) [15, 16]. Abnormal behavior is often readily apparent when cattle walk, as lameness is closely associated with painful gait and foot disease. The automation primarily focuses on detecting deviations between normal and abnormal behavior, classifying them to make decisions and further analysis.
However, the impact of such development in livestock farming is enormous, it reduces the labor costs, helps to identify health issues immediately, and increases production efficiency. Therefore, researchers are continuously working on the innovative smart farming solution with image processing and CV.
Fuentes et al. [14] used two Hikvision surveillance cameras to collect video data for their study. The cameras captured video from two different angles and extracted frames based on animal activity. They then used these frames to create a dataset, which was employed to train a model that effectively monitored the cows and performed statistical analysis to capture behavioral changes. McDonagh et al. [8] introduced a neural network model for behavior prediction by monitoring 46 individual pregnant cows. They specially considered the 10 hours prior video data to calving and focused on seven different behaviors. The model showed successful classification of the behavior over 80% instances and was able to predict calving before several hours with 83% accuracy. Fernández-Carrión et al. [10] proposed a deep learning model for the early detection of disease in Eurasian wild boar by monitoring their motion. They found a negative correlation between the animals’ movement and fever caused by the disease; the movement of infected animals was significantly reduced compared to that of healthy ones. In another study, Qiao et al. [15] proposed a combined model of Convolutional 3D and Convolutional Long Short-Term Memory (LSTM), which achieved higher accuracy compared to other state-of-the-art models in classifying five behaviors. Maria et al. [17]. focused on swine health, specifically identifying respiratory diseases. They considered eye temperature, heart rate, and respiration rate by extracting information from thermal infrared images. Their results demonstrated the successful implementation of CV in the early detection of respiratory diseases by considering physiological changes in pigs.
In another article, Barney et al. [19]. proposed a modified Mask-RCNN model for real-time automated lameness detection with high accuracy. In [14], the authors used Mask-RCNN and YOLO, both object detection algorithms. After detecting the cattle, they assigned a local ID to track each animal, then extracted features, and applied machine learning (ML) algorithms. Similarly, Russello et al. [5] presented a deep learning model to detect lameness by extracting nine annotations from video data. Wu et al. [18] introduced an LSTM model for large-scale farming to identify lame cows automatically. They used 50 videos, with a lame-to-non-lame sample ratio of 20:30. From these videos, they randomly selected 5,000 frames for their experiment, focusing solely on the position and movement of the cows’ legs. Their LSTM-based classification model achieved 98.57% accuracy, outperforming three traditional machine learning algorithms using a 50:50 train-test split.
However, the literature shows that cattle lameness was investigated in previous studies using surveillance video data collected from farms, with cameras installed at various angles. Additionally, most of those studies focused on small samples, and the investigations were conducted within limited and specific contexts. This study aimed to investigate lameness detection in large cattle farms, where hundreds of cattle need to be managed simultaneously. The research was conducted in two parts: video data collection and the development of a classification model. To achieve this, YouTube videos containing both lame and non-lame cattle were used. The labeling of normal and lame cattle was performed based on the title and description of each video. Since online videos were captured from diverse environments and locations, dataset diversity and robustness were ensured. To enhance the dataset size, vertical augmentation was applied to the training data. Moreover, to build the classification model, two deep learning classifiers, ConvLSTM2D and 3D CNN were utilized. The models were evaluated using multiple performance metrics. The classification results indicate successful implementation of online data with the proposed models, achieving comparable or superior accuracy compared to state-of-the-art models.
The remainder of this paper is organized as follows: Section II presents the materials and methods, including data collection, preprocessing, and model construction. Section III discusses the results obtained from the proposed models. Section IV discusses the findings and compares them with existing studies, along with limitations and future directions. Finally, Section V concludes the paper.
II Materials and Methods
Figure 1 illustrates the deep learning-based cattle lameness detection pipeline. The process begins with video data, from which frames are extracted during the data preprocessing stage. These extracted frames undergo data augmentation to enhance model generalization, producing a diverse training dataset. The model construction phase involves training two deep learning architectures: 3D CNN and ConvLSTM2D, denoted as $f_{\text{CNN3D}}(X)$ and $f_{\text{ConvLSTM2D}}(X)$, respectively. Finally, the models are evaluated using test data in the cattle classification stage to assess their performance.
II-A Dataset
The data used in previous relevant research papers were collected following a similar approach. First, a specific cattle farm, including cows, was selected. Then, different types of video cameras were installed in convenient locations on the grid. The cameras captured videos of the cattle from desired angles (e.g., walking, eating, and moving) for a specific amount of time. These videos were processed, and only the relevant parts were used for further analysis. Additionally, many studies performed locomotion scoring (a manual lameness detection approach) to classify cattle as lame or non-lame. Based on this classification, unique identification was assigned to the cattle, and tracking of both classes was maintained. Later, deep learning models were trained and assessed using the data.
II-A1 Data collection
The data collection process for this study was quite different from that of previous studies. We used an online platform to collect freely accessible cattle walking videos, specifically from YouTube. Publicly available data are commonly used in many research fields; for example, one of our previous articles [21] demonstrated that YouTube videos are frequently used in deepfake video detection research. In order to create an inclusive classification model, we created a diverse dataset. Therefore, the following features were incorporated with the video: there were different cattle in different environments, including intensive, semi-intensive, and extensive cattle farming; the cattle walked in different directions, including left to right and right to left; and the videos were captured from various views, including front, side, and back.
II-A2 Basic Information
TABLE I: Summary of the dataset statistics.
Class Video Count Duration (s) Frame Count
Normal 25 134.46 4050
Lame 25 148.91 5900
Total 50 283.37 9950
To create the dataset, a total of 50 videos featuring 42 cattle were considered, with an average duration of 5.67 seconds per video. We ensured an equal class distribution of the data, meaning that 25 lame and 25 nomal cattle videos were included. Table I presents the dataset distribution, showing the number of videos, total duration in seconds, and the corresponding frame counts for the normal and lame classes. The normal class videos have a cumulative duration of 134.46 seconds and consist of 4050 frames, whereas the lame class videos have a total duration of 148.91 seconds and comprise 5900 frames. Overall, the dataset includes 50 videos with a combined duration of 283.37 seconds and a total of 9950 frames. Besides, Figure LABEL:fig:all_images displays example frames from cattle video sequences, with upper three images representing lame cattle and the other representing normal cattle. The original videos had varying resolutions due to collection from different online sources. To standardize the input data, all videos were resized to a resolution of $500\times 500$ pixels. The dataset used in this study is available at the following Google Drive link: https://drive.google.com/drive/folders/1JHutEYjR6tAvXNnJuMhVgJabNC8tq7z7?usp=sharing.
Algorithm 1 Pseudo-code of the Proposed Classification Model
0:
  • $\bullet$
[…]
  • $\bullet$
    $N_{\text{frames}}$: Number of frames per video
  • $\bullet$
    $S_{\text{target}}$: Target size for frame resizing
  • $\bullet$
    $p_{\text{aug}}$: Data augmentation probability
  • $\bullet$
    Batch size $=B$, Number of epochs $=E$
0:  Trained 3DCNN or ConvLSTM2D model; performance metrics
1:  1. Data Preparation
2:  for $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{test}}$ do
[…]
4:     if $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}$ then
5:        $(X,y)\leftarrow(X\cup\{F_{i}\},y\cup\{l_{i}\})$
6:     else
7:        $(X_{\text{test}},y_{\text{test}})\leftarrow(X_{\text{test}}\cup\{F_{i}\},y_{\text{test}}\cup\{l_{i}\})$
8:     end if
9:  end for
10:  Apply horizontal flip augmentation to $X_{\text{train}}$ with probability $p_{\text{aug}}$
11:  Normalize $X_{\text{train}},X_{\text{test}}$ and one-hot encode labels
12:  2. Model Training
13:  $M\leftarrow\text{CreateModel}(N_{\text{frames}},S_{\text{target}})$
14:  $M\leftarrow\text{TrainModel}(M,X_{\text{train}},y_{\text{train}},B,E)$
15:  3. Model Evaluation
16:  $\hat{y}\leftarrow M(X_{\text{test}})$
17:  metrics,confusion $\text{metrics},\text{confusion matrix}\leftarrow\text{Evaluate}(y_{\text{test}},\hat{y})$
18:  return  $M$, metrics, confusion matrix
II-B Data used in this study
TABLE II: Basic information of the dataset used in this study compared with the most relevant datasets in the literature
Study Train Frames Test Frames Test Environment Focused Activity
1 1376 345 Within the barn Intensive farming
2 1000 500 Passing alley Intensive farming
3 490 210 Walked along a narrow path Intensive farming
4 690 341 Barn to milking parlor pathway Intensive farming
5 600 120 Corridor after the milking parlor Intensive farming
This study 1500 500 Indoor and outdoor environment Semi-intensive farming
Table II provides an overview of the data used in this research alongside datasets from previous cattle lameness detection studies. The table compares these studies based on the number of training and testing frames, the test environment, and the focused activity. As previously mentioned, the dataset consists of a total of 50 videos, including 25 normal and 25 lame cattle videos. Out of these, 30 videos (15 normal and 15 lame) were used for training the model, while the remaining 20 videos (10 normal and 10 lame) were used for testing. Previous studies focused on intensive farming environments, where cattle were primarily managed in indoor grazing conditions and data was collected from specific, limited locations such as within a barn, a passing alley, a narrow path, and corridors related to milking parlors. In contrast, this study focuses on a semi-intensive farming setting and incorporates both indoor and outdoor environments. This approach captures a wider range of cattle movement behaviors, lighting conditions, and surface variations, potentially leading to a more generalizable model for cattle lameness detection.
II-B1 Data pre-processing
Preprocessing in this investigation was conducted in several distinct steps, including frame extraction and data augmentation. The overall average frame rate was approximately 35.12 frames per second. However, due to the high volume of frames, only 25 images were randomly selected from each training video, resulting in a total of 750 images. These images were then cropped to a 224×224 pixel ratio. However, data augmentation is a widely recognized method for addressing limited data issues and helps to create robust and sustainable classification models. Numerous data augmentation techniques are available, as demonstrated by Jia et al. [3], who employed nine different methods, such as Mixup, Cutout, CutMix, etc. This study considered only image flipping as an augmentation technique, applied solely to the training data. All 30 training videos (750 frames) were virtually flipped, generating a total of 1500 images. In test data, 500 images were extracted from 20 videos. However, no data augmentation was performed on the test dataset to preserve integrity and ensure a fair assessment of the model’s generalization ability.
II-C Model Construction
The algorithm 1 outlines the pseudocode of the classfication model. It begins with data preparation by processing the training and testing datasets ($\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{test}}$), extracting frames from each video, resizing them to a target size ($S_{\text{target}}$), and ensuring each video has $N_{\text{frames}}$ frames through padding or truncation, creating input data ($X$) and labels ($y$). Training data is augmented with random horizontal flipping (probability $p_{\text{aug}}$). Pixel values are normalized, and behavior labels (“normal” and “lame”) are encoded numerically. Next, model training involves creating a 3D CNN or ConvLSTM2D model, configured using $N_{\text{frames}}$ and $S_{\text{target}}$, and training it on the prepared data ($X_{\text{train}}$, $y_{\text{train}}$) using a batch size ($B$) and epochs ($E$). The model learns patterns indicative of different behavior classes. Finally, model evaluation uses the trained 3D CNN model to predict behavior labels ($\hat{y}$) for the testing dataset ($X_{\text{test}}$), comparing them to true labels ($y_{\text{test}}$) to assess performance. Metrics like accuracy are calculated, and a confusion matrix is generated. The trained model, performance metrics, and confusion matrix are returned as output.
II-C1 3D CNN
CNNs are widely used deep learning models for image processing tasks. 3D CNNs are an extension of these and can handle 3D data containing spatiotemporal information, such as videos. They utilize 3D convolutional kernels, which enable them to focus on spatial features extracted from input data and distinguish themselves from traditional 2D CNNs. Because of their ability to learn both spatial and temporal features, this study employed a 3D CNN model using the Keras Sequential API to perform binary classification, where movement patterns are assessed over time. The network accepts input tensors of shape $(25,224,224,3)$, where 25 denotes the number of consecutive frames, $224\times 224$ represents the spatial resolution of each frame, and 3 corresponds to the RGB color channels.
As the Figure 2 shows, the model architecture begins with a 3D convolutional layer comprising 32 filters with a kernel size of $(3\times 3\times 3)$, employing a ReLU activation function and same padding to preserve the original spatial and temporal dimensions. This is followed by a 3D max pooling layer with a pooling size of $(2\times 2\times 2)$ to reduce dimensionality and computational load. A second convolutional layer with 64 filters of size $(3\times 3\times 3)$ and ReLU activation is subsequently applied, again followed by a $(2\times 2\times 2)$ max pooling operation. The extracted feature maps are then flattened into a one-dimensional vector, which is passed through two fully connected layers with 128 and 64 neurons, respectively, each utilizing a ReLU activation function. To reduce the risk of overfitting, dropout layers with a dropout rate of 0.5 are inserted after each dense layer. The final output layer consists of a single neuron with a sigmoid activation function, yielding a probability score for binary classification. The model is compiled using the Adam optimizer, binary cross-entropy as the loss function, and accuracy as the evaluation metric.
Figure 2: Architectures of the 3D CNN and ConvLSTM2D models used for the cattle classification.
II-C2 ConvLSTM2D
ConvLSTM2D is a deep learning architecture that integrates the strengths of spatial feature extraction and temporal dependency modeling through the combination of CNNs and LSTM networks, respectively. This makes it particularly effective for understanding patterns and movements in sequential data. Followed by the previous model, the input layer is similar, processing video frames in the same format. In Figure 2, model 2 employed in this study begins with a ConvLSTM2D layer. Following the ConvLSTM2D layer, two successive 3D max pooling layers with pooling sizes of $(1\times 2\times 2)$ are applied to downsample the spatial dimensions while maintaining the temporal dimension. The output is then flattened into a one-dimensional feature vector. A fully connected dense layer with 128 neurons and ReLU activation is employed for high-level feature abstraction, followed by a dropout layer with a rate of 0.25 to reduce overfitting. Finally, a dense layer with sigmoid activation is used to classify cattle into ‘normal’ and ‘lame’ categories. The compilation process also follows the same procedure as in the previous model.
Figure 3: Confusion matrix of both models under the same test data.
II-D Model Evaluation
The four performance evaluation measures (accuracy, precision, recall, and F1-score) used in this study are associated with the confusion matrix; therefore, both are presented gradually. The values of all evaluation criteria range from 0% to 100%, where a higher value indicates stronger performance. In the mathematical framework for model evaluation, $n$ represents the total instances, $y_{i}$ denotes the true label for instance $i$, and $\hat{y}_{i}$ signifies the model’s predicted label. The indicator function $I(\text{condition})$ returns 1 if the condition is true, otherwise 0. Summation, denoted by $\sum_{i=1}^{n}$, calculates the cumulative value across all instances. These components enable quantitative assessment by comparing predicted ($\hat{y}_{i}$) and true ($y_{i}$) labels, facilitating calculation of the evaluation metrics, providing a comprehensive evaluation of model performance.
$Accuracy=\frac{1}{n}\sum_{i=1}^{n}I(y_{i}=\hat{y}_{i})$ (1)
$Precision=\frac{\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=1)}{\sum_{i=1}^{n}I(\hat{y}_{i}=1)}$ (2)
$Recall=\frac{\sum_{i=1}^{n}I(y_{i}=1\land\hat{y}_{i}=1)}{\sum_{i=1}^{n}I(y_{i}=1)}$ (3)
$F1-Score=\frac{2\times Precision\times Recall}{Precision+Recall}$ (4)
III Results
The dataset utilized in this study comprised 50 videos, with 50% depicting defective cattle and 50% depicting non-defective cattle. Of these 50 videos, 30 were used for model training, and 20 were used for validating the 3D CNN and ConvLSTM2D models. This project was developed using the Python programming language and implemented within the Google Colab Pro environment. Data preprocessing tasks were handled by OpenCV and MoviePy, supplemented by Keras’ ImageDataGenerator for data augmentation. Furthermore, TensorFlow, NumPy, Pandas, Scikit-learn, and Matplotlib were employed for model construction, training, performance evaluation, and visualization. The classification results achieved by the models were calculated and presented below.
The models generate predictions for each individual frame within each video in the test data. It is assumed that if the model correctly predicts the majority of frames in a video, then the entire video is considered correctly classified, despite the presence of a few incorrect frame-level predictions. The confusion matrices for both models are presented in Figure 3. As shown in Figures 3(a) and 3(b), a small number of misclassifications were observed, including 1 false positives and 1 false negative for 3D CNN, and 1 false positives and 2 false negatives for ConvLSTM2D.
The 3D CNN model demonstrated notable performance in classifying the cattle data. Table III indicates an accuracy of 90%, along with strong average precision, recall, and F1-scores. On the other hand, ConvLSTM2D achieved comparatively lower performance, with accuracy, precision, recall, and F1-score values of 85%, 90%, 81.82%, and 85.71%, respectively. Overall, the confusion matrices and classification reports highlight the models’ effectiveness in classifying healthy and defective cattle.
TABLE III: Lameness classification results of proposed 3D CNN and ConvLSTM2D
Model Accuracy Precision Recall F1-score
3DCNN 90 90.9 90.9 90.91
ConvLSTM2D 85 90 81.82 85.71
IV Discussion
TABLE IV: Performance comparison of this study with existing cattle lameness classification models
Study Classifier Accuracy Precision Recall F1-score
[4] Logistic Regression 81.18 75 47.06 57.83
[15] C3D-ConvLSTM 90.32 90.62 88.76 89.68
[5] SVM 80.07 78.70
[20] AdaBoost 77.9
[22] DT 90.18
[23] CNN 90.21
[24] SVM 83.58 86.04 88.43 87.21
[25] KNN 73.33 89.08 75.28 81.60
This study 3DCNN 90 90.9 90.9 90.91
  • – No specific data was found
Based on available information, all previous research on cattle lameness, sickness, or gait abnormality detection using computer vision technology has been conducted with very specific and limited data types. Typically, these studies focused on a single cattle barn, where video cameras were set up to monitor the cattle. The collected video data were then used for further analysis, including object detection, pose estimation, feature extraction, and classification. In contrast, this study utilized cattle videos from online platforms, which encompass multi-dimensional features. This departure from previous studies allowed for the development of a more universal and comprehensive lameness detection model that is suitable for large-scale cattle farms. The results of this study also validate the applicability of the classification model.
Table IV presents a performance comparison between eight prior studies and the proposed model. The table evaluates models based on accuracy, precision, recall, and f1-score. The final row of the table displays the results of this study, which employed a 3D CNN model, while the preceding rows summarize the findings from other relevant studies that utilized various machine learning and deep learning classifiers. The model used in this study achieved a notable accuracy of 90%. More significantly, it demonstrated the highest precision among the studies where precision values were available, reaching 90.9%. This indicates that the model exhibits a strong ability to minimize false positives, correctly identifying positive cases with high confidence. Additionally, the 3D CNN model also yielded a robust F1-score of 90.91%. When compared to other approaches, the 3D CNN model in this study demonstrated competitive performance, whereas study [15] employed C3D-ConvLSTM achieved a slightly higher accuracy of 90.32% but a lower precision of 90.62%.
The results indicate that deep learning models generally outperform traditional machine learning approaches, such as SVM, Logistic Regression, and KNN. Furthermore, the findings of this study demonstrate the effectiveness of 3D CNNs for the video classification, as they efficiently capture both spatial and temporal information and offer well-balanced performance across multiple evaluation metrics.
IV-A Limitation and future plan
While this study demonstrates the potential of deep learning for cattle lameness detection, certain limitations may affect its broader applicability and real-world deployment, potentially impacting the generalizability of the findings to broader contexts. The following points outline these limitations:
  • All of the videos were sourced from YouTube. The cattle abnormality was determined based on the video titles and descriptions; however, it was not possible to individually verify the class and level of abnormality.
  • Although the simplicity of the proposed methodology (which directly employs classification algorithms) enables successful implementation with the video data, it is also necessary to experiment with intermediate processing steps such as object detection and pose estimation.
  • While this study utilized a greater number of training and testing frames compared to many similar studies, the total number of video samples remains relatively limited, which may affect the model’s generalizability and robustness.
To address the limitations outlined in this study and further enhance the applicability of our approach, future work should focus on expanding the dataset with a greater variety of video sources, including direct recordings from cattle farms, which would ensure a more robust representation of real-world scenarios, encompassing diverse cattle breeds, environmental conditions, and lameness severities, as well as addressing potential biases in online video content. In addition, while this study initially utilizes a direct classifier, future research will explore the impact of feature engineering through object detection, pose estimation, and feature extraction to assess the dataset and model performance. Overall, this study is an initial initiative of a benchmark cattle lameness detection dataset, future work will be introducing a complete benchmark of the approach.
V Conclusions
This paper presents a cattle lameness video dataset and deep learning-based methods for effectively distinguishing between lame and non-lame cattle from video data. Compared to existing datasets used for cattle lameness detection, this study collected data entirely from online sources, providing more comprehensive information across diverse environments, scenes, and activities. Additionally, two deep learning models, 3D CNN and ConvLSTM2D, were used to classify normal and lame cattle. From the dataset, 1,500 images were extracted for training the models, while 500 images were used for validation. The experimental results demonstrated improved or at least comparable performance to existing cattle lameness classification models. In conclusion, this study serves as a reference for detecting cattle lameness or behavior without relying on traditional multi-step approaches.